# but Need a Good Lens

### v1

@@ Appendix D Statistical Feasibility on Other LLMs
As shown in Table 4 and 5, meta-cognition lenses perform weakly in capturing meta-cognition of Llama-3-8B-Instruct and Mistral-7B-Instruct. Interestingly, however, these methods are statistically better at hard tasks than simple ones. For instance,
@@
CoE-C 0.179 (8.0e-10) 0.119 (1.1e-9) 0.095 (4.0e-8) 0.054 (3.1e-6) 0.203 (8.2e-15) 0.135 (2.0e-14)
CoE-R 0.219 (3.8e-14) 0.139 (1.4e-12) 0.123 (1.2e-12) 0.083 (9.2e-13) 0.155 (3.8e-9) 0.103 (4.5e-9)
Maxprob 0.102 (0.0005) 0.081 (3.6e-5) 0.098 (1.5e-8) 0.063 (6.9e-8) 0.022 (0.4058) 0.013 (0.4569)
PPL 0.100 (0.0006) 0.080 (4.2e-5) 0.101 (6.1e-9) 0.064 (3.4e-8) 0.263 (2.2e-22) 0.180 (9.1e-23)
Entropy 0.119 (5.0e-5) 0.092 (2.3e-6) 0.104 (2.0e-9) 0.065 (2.6e-8) 0.044 (0.098) 0.027 (0.127)
$\Delta$Entropy -0.208 (8.0e-13) -0.136 (3.4e-12) 0.024 (0.167) 0.019 (0.094) -0.074 (0.0047) -0.049 ( 0.00499)
@@
CoE-C 0.194 (9.0e-08) 0.130 (1.1e-07) 0.215 (3.9e-17) 0.139 (5.9e-16) 0.333 (2.5e-18) 0.221 (2.7e-17)
CoE-R -0.282 (3.9e-15) -0.190 (9.1e-15) -0.118 (4.8e-06) -0.078 (5.8e-06) -0.101 (0.0099) -0.064 (0.0148)
Maxprob 0.306 (1.1e-17) 0.203 (1.1e-16) 0.036 (0.158) 0.023 (0.185) 0.011 (0.783) 0.005 (0.848)
PPL 0.320 (3.1e-19) 0.213 (3.7e-18) 0.051 (0.048) 0.032 (0.061) 0.033 (0.396) 0.020 (0.437)
Entropy 0.326 (6.8e-20) 0.217 (7.2e-19) 0.035 (0.181) 0.023 (0.191) 0.012 (0.754) 0.007 (0.794)
$\Delta$Entropy 0.154 (2.4e-05) 0.105 (2.1e-05) 0.009 (0.713) 0.006 (0.721) -0.040 (0.303) -0.027 (0.301)
CoE-C has a stronger correlation with PRM rewards on MinervaMATH than GSM8k and MATH500 for the two LLMs.

### v2

@@ Appendix D Statistical Feasibility on Other LLMs
As shown in Table 6 and 7, meta-cognition lenses perform weakly in capturing meta-cognition of Llama-3-8B-Instruct and Mistral-7B-Instruct. Interestingly, however, these methods are statistically better at hard tasks than simple ones. For instance, CoE-C has a stronger correlation with PRM rewards on MinervaMATH than GSM8K and MATH500 for the two LLMs.
@@
CoE-C
0.179
(8.0e-10)
0.119
(1.1e-9)
0.095
(4.0e-8)
0.054
(3.1e-6)
0.203
(8.2e-15)
0.135
(2.0e-14)
CoE-R
0.219
(3.8e-14)
0.139
(1.4e-12)
0.123
(1.2e-12)
0.083
(9.2e-13)
0.155
(3.8e-9)
0.103
(4.5e-9)
Maxprob
0.102
0.081
(3.6e-5)
0.098
(1.5e-8)
0.063
(6.9e-8)
0.022
0.013
PPL
0.100
0.080
(4.2e-5)
0.101
(6.1e-9)
0.064
(3.4e-8)
0.263
(2.2e-22)
0.180
(9.1e-23)
Entropy
0.119
(5.0e-5)
0.092
(2.3e-6)
0.104
(2.0e-9)
0.065
(2.6e-8)
0.044
0.027
[…]
-0.208
(8.0e-13)
-0.136
(3.4e-12)
0.024
0.019
-0.074
-0.049
( 0.00499)
@@
CoE-C
0.194
(9.0e-08)
0.130
(1.1e-07)
0.215
(3.9e-17)
0.139
(5.9e-16)
0.333
(2.5e-18)
0.221
(2.7e-17)
CoE-R
-0.282
(3.9e-15)
-0.190
(9.1e-15)
-0.118
(4.8e-06)
-0.078
(5.8e-06)
-0.101
-0.064
Maxprob
0.306
(1.1e-17)
0.203
(1.1e-16)
0.036
0.023
0.011
0.005
PPL
0.320
(3.1e-19)
0.213
(3.7e-18)
0.051
0.032
0.033
0.020
Entropy
0.326
(6.8e-20)
0.217
(7.2e-19)
0.035
0.023
0.012
0.007
[…]
0.154
(2.4e-05)
0.105
(2.1e-05)
0.009
0.006
-0.040
-0.027